# A Cloud-native Agile approach to cyber platform prototyping and integration for astronomy: the ENGAGE SKA case

### v1

@@ 2 State of the Art
  • Change management: ie Subversion[11]
@@
The Container Orchestration at SKAO supporting cloud infrastructure follows a common containerization framework with portability as a key attribute with key items: Resource abstraction; Scheduling & Deployment; Scaling & Load balancing; Monitoring, Auto-healing & resource allocation management. A container is essentially a fully packaged and portable computing environment and as such it allows great flexibility while encapsulating and isolating the application with everything an application needs for proper execution: binaries, libraries, configuration files, and dependencies. The container itself is abstracted away from the host OS, with only limited access to underlying resources—much like a lightweight virtual machine (VM). As a result, the containerized application can be run on various types of infrastructure—on bare metal, within VMs, and in the cloud without needing to refactor it for each environment. Container Orchestration is the mechanism by which large scale deployment of containers and dependent resources (network,storage,GPU, etc.) is managed and scaled across a multi-node and multi-data centre managed infrastructure environment. This is done seamlessly so that the combined infrastructure can be considered a single resource. The Cloud Native Monitoring & Control layer follows the following description:
@@
On the other hand, the production and exploitation of industrial control systems that are applicable to the thousands of devices that will operate and monitor SKAO differs substantially from traditional information systems; this is in part due to constraints on the availability and changes in the life cycle of production systems, as well as their reliance on proprietary protocols and software packages with little support for open development standards. The application of agile software development methods therefore represents a challenge which requires the adoption of existing change and build management tools and approaches that can help bridge the gap and reap the benefits of managed development when dealing with industrial control systems. For instance, agile development tools such as Apache Maven for build management, Hudson for continuous integration or Sonatype Nexus for the operation of "definite media libraries" were leveraged to manage the development lifecyle of systems similar to the CERN UAB framework [14], a major large scale research infrastructure requiring advanced compute platforms for its sensors’ network. This software (that became by itself a well studied standard and a case study influencing many other projects in high energy physics - accelerators and synchrotrons - and radioastronomy) became included the CERN Common Middleware or the Front-end Software Architecture (FESA) project[15].
The SKAO has its observatory software tree structured in several Agile Array Release Trains (ARTs), namely the Observation Management and Control (OMC), the Data processor (DP) ART, the Services ART, MID Integration and Low Integration ARTs. Here, the OMC is meant to address proposal submission & management, design, scheduling, and execution of observations on both MID and LOW telescopes; correctly operate and monitor the observatory and the telescopes (including Local Monitoring and Control and the Central Signal Processor Local Monitoring and Control ); guide the development of TANGO device interfaces; guide the Integration of engineering operations’ software products.
The Services ART is meant to provide and develop the Compute Platforms, the Networks and the software services to the other ARTs, namely the OMC, MID and LOW Integration. These include: Elasticstack search infrastructure, deployment and management of the Observatory Science Operations Data Archive (ODA), the Observatory Engineering Data Archive (EDA), CI/CD infrastructure (built on GitLab), Jira, Confluence, SKAO email support to Slack, Miro, Google, Platform monitoring, Developer reference implementations, tooling and documentation.
@@
Within the Service ART, ENGAGE SKA has supported the System team, a specialized Agile Team that assists in building and using the Agile development environment, including CI/CD and test automation. The System Team supports the integration of assets from SKAO Agile teams, performs end-to-end solution testing when necessary, and assists with deployment and release. The SKAO implemented one single System Team right before construction, from the bridging phase. This team is responsible for setting the basis and enable the project wide automations needed to implement CI/CD and DevOps. This work is based on the harmonization of previous experiences and practices, as emerged in the pre-construction phase from the experience developed by different consortia. Up to and during construction, the team also contributes to tasks related to system Critical Design Review (CDR) artefacts refinement.
@@
Over the years, major astronomical observatories like the Atacama Large Millimeter / submillimeter Array (ALMA)[16], the biggest operating millimeter wave interferometer to date, the Vera Rubin Observatory[17] soon to become the largest wide sky survey machine in the optical domain, the LOw Frequency ARray (LOFAR)[18] and the Giant Metrewave Radio Telescope (GMRT)[19] currently the two largest long radio wave interferometers in operation), and the four SKA precursors : MeerKAT[20] and the Hydrogen Epoch of Reionization Array (HERA)[21], in South Africa, and the two radio telescopes in Western Australia, the Australian SKA Pathfinder (ASKAP)[22] and the Murchison Widefield Array (MWA)[23], besides many other large scale observatories, have developed their own customized solutions for their software lifecycle tackling a variety of conditions imposed by their sheer size and often by the nature of their multinational and distributed collaborations. Although SKAO has chosen state of the art trends from the Cloud native world, the ART teams took into account a variety of lessons learned from these projects on the design and decision making process associated with cyberinfrastructure and software lifecycles. We briefly describe in the following subsections the major characteristics of ALMA and Vera Rubin.
2.1 ALMA Observatory
The Atacama Large Millimeter/submillimeter Array (ALMA) was built and is operated through a joint effort from partners on three different continents: the European Southern Observatory (ESO), the U.S. National Science Foundation (NSF) and the National Institutes of Natural Sciences (NINS) of Japan. ALMA’s Integrated Computing Team (ICT) coordinates the ALMA software engineers working across the different regional groups within the the Joint ALMA Observatory (JAO). The software development and build procedures at ALMA followed a more monolithic approach. The ALMA Observatory software counts with roughly 11 millions lines of code, runs on about 350 hosts and is developed at 7 sites around the globe.The ALMA Software build strategy was focused on abstraction of compilation and linking procedures for non-programmers, but forwent inter-dependency between subsystems, and their modules, as described in [24]. During its phased construction and deployment it gradually added capabilities as required by the scientific community to enhance science productivity and it moved towards Continuous Integration and deployment as the chosen software development practices, cutting down compilation times down from 24 hours to 5 hours, and enabling efficient bug tracking and correction by ASO teams during compilation and deployment. Regularly, Yearly Observing Cycle several software releases were prepared, incorporating new functionalities.
The ALMA software configuration control uses a Subversion (SVN) under the Apache License as a Versioning Control System (VCS). The repository has a tree-structure reflecting the various subsystems with their own interdependencies. These implies a careful makefile configuration to build and test the entire ALMA software. The software is developed and deployed on customized Red Hat Enterprise Linux systems, and relies for most of the subsystems on the in-house developed ALMA Common Software (ACS) as middleware/framework [25]. Jenkins tools were used at ESO[26] as a continuous integration tool, to provide services such as code builds, code tests, calculation of code metrics and checks of coding standards while Atlassian’s JIRA tool was used for issue tracking. Being a partnership between three different regions, the deployment of accepted software is done simultaneously at the JAO and the different ALMA Regional Centers (ARCs). The management of the ever-growing observational data makes extensive usage of a virtualized environment with a container-based database allowing the creation of a full copy of the production database in a very short time[27].
2.2 Vera Rubin Observatory
The Vera Rubin Observatory, is an astronomical observatory currently under construction in Chile. Its main task will be carrying out a synoptic astronomical survey, the Legacy Survey of Space and Time (LSST). The Rubin Observatory will house the Simonyi Survey Telescope[17], a wide-field reflecting telescope with an 8.4-meter primary mirror[28, 29] that will photograph the entire available sky every few nights.
The LSST data management system (LSSt DMS) will be a world-class computational facility that will control the telescope, process and deliver LSST data. The total amount of data collected over the ten years of LSST operation will top 60 petabytes (PB) and producing a 20 PB catalog database. To enable this processing, the LSST project is developing a new, general-purpose, high-performance, scalable, well documented, open source data processing software stack for Optical and Infra Red surveys. Prototypes of this stack are already capable of processing data from existing cameras (e.g., SDSS, DECam, MegaCam), and form the basis of the Hyper-Suprime Cam (HSC) Survey data reduction pipeline[30]. The total data volume after processing will be several hundred PB, processed using about 150 TFLOPS of computing power for the first Data Release, increasing to 950 TFLOPS by Data Release 11 at the end of the ten-year survey. The computational facilities and data archives of the LSST DMS will become one of the largest and most important science facilities of its kind in the world.
LSST software relies on Phalanx[31], a GitOps repository for Rubin Observatory’s Kubernetes environments, notably including Rubin Science Platform deployments[32]. Using Helm and Argo CD, Phalanx defines the configurations of applications in each environment. Phalanx is a SQuaRE deployment (Science Quality and Reliability Engineering, the team responsible for the Rubin Science Platform). This set of tools provides documentation for Rubin team members that are developing applications and administering Kubernetes (K8s)[33] clusters that are best suited to respond to the challenges of running a service-oriented architecture in a modern observatory environment. Furthermore, a whole suite of software for Data management and access to several levels of data is available to Astronomers and other end-users via the Rubin Documentation Portal to learn how to use Rubin Observatory’s software, services, and datasets.
@@ 3 The SKAO software integration environment
At this stage, the artefacts delivered should be seen as fully fledged systems (software running on integrated platform servers) capable of providing end-to-end value to the multiple stakeholders. All the tests that are included with SKAMPI were therefore executed with the objective of verifying and validating end-to-end capability. How those artefacts are assembled, configured and tested is determined by the source code defined within the SKAMPI repository. The software included in every version of SKAMPI consists of a software base with common parts, resulting in the following deliverable artifacts:
  • SKA Mid software component;
  • SKA Low software component;
  • SKA Common/HQ software component;
@@ 3.2 The SKAMPI Infrastructure
As such, SKAMPI was deprecated in favour of a system with better ownership definitions that allowed for integration between two distinct components of OMC instead of all of them. SKAMPI was a necessary learning path for software integration that enabled through its "lessons learned" a substantial and important experience towards a functional OMC.
@@ 4.1 Tango
The SKAO has selected TANGO-controls as the distributed middleware for driving software and hardware devices using the Common Object Request Broker Architecture (CORBA)[41] distributed objects that represent devices that communicate with ZeroMQ[42] events internally. The entire system runs on a containerised environment managed by Kubernetes (k8s) with Helm[43] for packaging and deploying SKA software. As described in [39], k8s uses a declarative “model” of operation that drives the system towards the desired state described by user manifests, with various controller components managing the lifecycle of the associated resources. Helm provides the concept of a chart which is a recipe to deploy multiple k8s resources (i.e., containers, storage, networking components, etc…) required for an application to run. The resources are created using templates so that the chart can adapt generic configurations to different environments (i.e., the different SKA datacentres). The SKA deployment practices include the heavy use of standardised Makefiles (i.e., for building container images, for testing, for the deployment of a chart, etc.) and Gitlab for the CICD (continuous integration continuous deployment) practices.
@@
4.2 Webjive / Taranta
Taranta is a Webjive based web application that allows a user to create a graphical user interface to interact with Tango devices. The interface may include a variety of charts, numerical indicators, dials, commands that can be used to monitor and to control devices. Very little knowledge of web technologies is needed. Webjive under the SKA organization will be used to control and monitor hundreds of antennas and telescopes spread across different areas. In the current phase of SKA process Webjive is being used as a tool for SKA developers to debug and test their simulated tango devices. Webjive is the graphical interfaced adopted by almost all teams present in the SKA project, spread across all continents and taking advantage of the cloud current setup implemented in SKA; this allows developers from different teams, co-located or geographically separated to access the same interface and the same functionalities available to control/debug their or other teams devices.
WebJive/Taranta is composed of four projects:
  • WebJive Frontend[44]
  • WebJive Dashboards[45]
  • WebJive Auth[46]
  • TangoGQL[47]
The Figure 6 diagram shows an overview of the Taranta architectue.
Figure 6: WebJive Architecture.
The Taranta Frontend is a React [^2] https://react.dev/ client that is used both to browse, inspect and control Tango devices and to create and run dashboards, each composed of widgets. React itself is a library for web and native user interfaces out of individual pieces called "components" that are JavaScript functions. These independent components can be seamlessly combined either individually or through the work of thousands of developers. Taranta Frontend accesses the Tango Control Framework through TangoGql, a GraphQL interface to Tango. The react-dom/client APIs let the developer render React components on the client (ie, in its browser). These APIs are typically used at the top level of an app to initialize a React tree. A customized framework may call the APIs through appropriate automation. React can be understood also as an architecture and its frameworks that implement it enable the user to fetch data in asynchronous components that run on the server or even during the build of the code, essentially turning the user/developer to be a web and a native developer.
The communication between Taranta and TangoGQL is managed by an appropriate frontend component. To use Taranta, an anonymous user can run dashboards and can browse and inspect devices. However, to be able to send commands to devices, to change their attributes and to create or modify dashboards, you need to authenticate with Taranta. Taranta uses taranta-auth to manage users; it accesses an Lightweight Directory Access Protocol (LDAP) repository or a JavaScript Object Notation (JSON) file to retrieve user information.
The dashboards created by an authenticated user are stored in a MongoDB database through the dashboard-repo application.
@@ 6 Conclusions
The current trends in the area of Green ICT, embodied in the Cloud native Computing technologies and the DevOps software ideas are influencing the design of compute infrastructures that will be key to reduce the operational costs of SKAO while improving the control and monitoring of the SKAO antennas and ancillary systems, Correlators, HPC facilities or related data centre tiered systems. Cloud native environment provides a target environment and framework for developers to aim for and enables a greater possible portability from desktop to production. Additionally, it insulates activity from production platform concerns with the latest possible practical hardware commitment, lowers barrier to entry, reduces time-to-share and enables integration and testing in a completely virtual environment by providing an abstraction from the underlying infrastructure.
@@ Acknowledgments
This research was supported by the project Enabling Green E-science for the SKA Research Infrastructure (ENGAGE SKA), reference POCI-01-0145-FEDER-022217, funded by COMPETE 2020 and FCT, Portugal and Project Centro de Investigação em Ciências Geo-Espaciais, reference UIDB/00190/2020, funded by COMPETE 2020 and FCT, Portugal; JPB acknowledges support from Project Lab. Associado UID/EEA/50008/2019. MdC has been supported by the Italian Government (MEF- Ministero dell’Economia e delle Finanze, MIUR - Ministero dell’Istruzione, dell’Università e della Ricerca). The ENGAGE SKA team acknowledges all the support and collaboration from the SKA Observatory.

### v2

@@ 2 State of the Art
  • Change management: ie Subversion (SVN)[11]
@@
Although SVN was initially adopted, the standardization of GitLAb made its way and this tool quickly became the open-source version control system and a cornerstone of SKAO DevOps change management. The Container Orchestration at SKAO supporting cloud infrastructure follows a common containerization framework with portability as a key attribute with key items: Resource abstraction; Scheduling & Deployment; Scaling & Load balancing; Monitoring, Auto-healing & resource allocation management. A container is essentially a fully packaged and portable computing environment and as such it allows great flexibility while encapsulating and isolating the application with everything an application needs for proper execution: binaries, libraries, configuration files, and dependencies. The container itself is abstracted away from the host OS, with only limited access to underlying resources—much like a lightweight virtual machine (VM). As a result, the containerized application can be run on various types of infrastructure—on bare metal, within VMs, and in the cloud without needing to refactor it for each environment. Container Orchestration is the mechanism by which large scale deployment of containers and dependent resources (network,storage,GPU, etc.) is managed and scaled across a multi-node and multi-data centre managed infrastructure environment. This is done seamlessly so that the combined infrastructure can be considered a single resource. The Cloud Native Monitoring & Control layer follows the following description:
@@
On the other hand, the production and exploitation of industrial control systems that are applicable to the thousands of devices that will operate and monitor SKAO differs substantially from traditional information systems; this is in part due to constraints on the availability and changes in the life cycle of production systems, as well as their reliance on proprietary protocols and software packages with little support for open development standards. The application of agile software development methods therefore represents a challenge which requires the adoption of existing change and build management tools and approaches that can help bridge the gap and reap the benefits of managed development when dealing with industrial control systems. For instance, agile development tools such as Apache Maven for build management, Hudson for continuous integration or Sonatype Nexus for the operation of "definite media libraries" were leveraged to manage the development lifecyle of systems similar to the CERN UAB framework [14, 15].
@@
The SKAO has its observatory software tree structured in several Agile Array Release Trains (ARTs), namely the Observation Management and Control (OMC), the Data processor (DP) ART, the Services ART, MID Integration and Low Integration ARTs. Here, the OMC is meant to address proposal submission & management, design, scheduling, and execution of observations on both MID and LOW telescopes; correctly operate and monitor the observatory and the telescopes (including Local Monitoring and Control and the Central Signal Processor Local Monitoring and Control ); guide the development of TANGO device interfaces; guide the Integration of engineering operations’ software products. The Services ART is meant to provide and develop the Compute Platforms, the Networks and the software services to the other ARTs, namely the OMC, MID and LOW Integration. These include: Elasticstack search infrastructure, deployment and management of the Observatory Science Operations Data Archive (ODA), the Observatory Engineering Data Archive (EDA), CI/CD infrastructure (built on GitLab), Jira, Confluence, SKAO email support to Slack, Miro, Google, Platform monitoring, Developer reference implementations, tooling and documentation. Within the Service ART, ENGAGE SKA has supported the System team, a specialized Agile Team that assists in building and using the Agile development environment, including CI/CD and test automation. The System Team supports the integration of assets from SKAO Agile teams, performs end-to-end solution testing when necessary, and assists with deployment and release. The SKAO implemented one single System Team right before construction, from the bridging phase. This team is responsible for setting the basis and enable the project wide automations needed to implement CI/CD and DevOps. This work is based on the harmonization of previous experiences and practices, as emerged in the pre-construction phase from the experience developed by different consortia. Up to and during construction, the team also contributes to tasks related to system Critical Design Review (CDR) artefacts refinement. The system Team acted in fact as an operational Agile analogue to a Change Control Board (CCB) for all the software modules version control by the other Agile teams in SKAO and articulated with the Feature and Product Owners of the several ART teams on software Quality Assurance matters.
@@
Over the years, major astronomical observatories like the Atacama Large Millimeter / submillimeter Array (ALMA)[16], the biggest operating millimeter wave interferometer to date, the Vera Rubin Observatory[17] soon to become the largest wide sky survey machine in the optical domain, the LOw Frequency ARray (LOFAR)[18] and the Giant Metrewave Radio Telescope (GMRT)[19] currently the two largest long radio wave interferometers in operation), and the four SKA precursors : MeerKAT[20] and the Hydrogen Epoch of Reionization Array (HERA)[21], in South Africa, and the two radio telescopes in Western Australia, the Australian SKA Pathfinder (ASKAP)[22] and the Murchison Widefield Array (MWA)[23], besides many other large scale observatories, have developed their own customized solutions for their software lifecycle tackling a variety of conditions imposed by their sheer size and often by the nature of their multinational and distributed collaborations. Although SKAO has chosen state of the art trends from the Cloud native world, the ART teams took into account a variety of lessons learned from these projects on the design and decision making process associated with cyberinfrastructure and software lifecycles.
@@ 3 The SKAO software integration environment
At this stage, the artefacts delivered should be seen as fully fledged systems (software running on integrated platform servers) capable of providing end-to-end value to the multiple stakeholders. All the tests that are included with SKAMPI were therefore executed with the objective of verifying and validating end-to-end capability. How those artefacts are assembled, configured and tested is determined by the source code defined within the SKAMPI repository. The software included in every version of SKAMPI consists of a software base with common parts, addressing the SKA Mid, SKA Low and SKA Common/HQ software components.
@@ 4.1 Tango
The SKAO has selected TANGO-controls as the distributed middleware for driving software and hardware devices using the Common Object Request Broker Architecture (CORBA)[41] distributed objects that represent devices that communicate with ZeroMQ[42] events internally. The entire system runs on a containerised environment managed by Kubernetes (K8s) with Helm[43] for packaging and deploying SKA software. As described in [39], K8s uses a declarative “model” of operation that drives the system towards the desired state described by user manifests, with various controller components managing the lifecycle of the associated resources. Helm provides the concept of a chart which is a recipe to deploy multiple K8s resources (i.e., containers, storage, networking components, etc…) required for an application to run. The resources are created using templates so that the chart can adapt generic configurations to different environments (i.e., the different SKA datacentres). The SKA deployment practices include the heavy use of standardised Makefiles (i.e., for building container images, for testing, for the deployment of a chart, etc.) and GitLab for the CICD (continuous integration continuous deployment) practices.
@@ 6 Conclusions
Cloud native environment provides a target environment and framework for developers to aim for and enables a greater possible portability from desktop to production. Additionally, it insulates activity from production platform concerns with the latest possible practical hardware commitment, lowers barrier to entry, reduces time-to-share and enables integration and testing in a completely virtual environment by providing an abstraction from the underlying infrastructure.
@@ Acknowledgments
This research was supported by the project Enabling Green E-science for the SKA Research Infrastructure (ENGAGE SKA), reference POCI-01-0145-FEDER-022217, funded by COMPETE 2020 and FCT, Portugal and Project Centro de Investigação em Ciências Geo-Espaciais, reference UIDB/00190/2020, funded by COMPETE 2020 and FCT, Portugal; JPB acknowledges support from Project Lab. Associado UID/EEA/50008/2019. MdC has been supported by the Italian Government (MEF- Ministero dell’Economia e delle Finanze, MIUR - Ministero dell’Istruzione, dell’Università e della Ricerca). The ENGAGE SKA team acknowledges all the support and collaboration from the SKA Observatory. DN acknowledges support from the HPC Chair at University of Évora.